Miahhhhhh:

Suppose a binomial event has a probability of success of 0.3, and 1250 trials are performed. What is the standard deviation of the possible outcomes? Round your answer to two decimal places

jhonyy9:

any idea ?

charge69:

The standard deviation for binomial distributions is: Sqrt(npq) N=trials=1250 P=p(success)=0.3 Q=p(failure)=1-p(success)=1-0.3=0.7 sqrt(1250×0.3×0.7) I think you can solve that part now.